# One-Shot Imitation under Mismatched Execution

### v2

@@ Abstract
Human demonstrations as prompts are a powerful way to program robots to do long-horizon manipulation tasks. However, translating these demonstrations into robot-executable actions presents significant challenges due to execution mismatches in movement styles and physical capabilities. Existing methods either depend on robot-demonstrator paired data, which is infeasible to scale, or rely too heavily on frame-level visual similarities that often break down in practice. To address these challenges, we propose RHyME, a novel framework that automatically aligns robot and demonstrator task executions using optimal transport costs. Given long-horizon robot demonstrations, RHyME synthesizes semantically equivalent demonstrator videos by retrieving and composing short-horizon demonstrator clips. This approach facilitates effective policy training without the need for paired data. We demonstrate that RHyME outperforms a range of baselines across cross-embodiment datasets, showing a 52% increase in task recall over prior cross-embodiment learning methods. We release our code and datasets at website.
@@ I Introduction
Human demonstrations offer an effective approach for programming robots to execute long-horizon manipulation tasks [1, 2, 3, 4]. Unlike language instructions, demonstrations are grounded in the task environment, providing rich with cues for what steps to follow, which objects to interact with, and how to interact with them [5, 6].
@@
We view this as a translation problem where a human video must be translated into a series of robot actions [7, 8, 9, 10]. However, training such policies typically requires paired human-robot demonstrations, which is impractical to collect at scale for long-horizon tasks. Although large-scale human demonstration videos (e.g., YouTube) and robot datasets exist [11, 12], they are unpaired, making them unsuitable for directly learning this translation.
@@
Prior works leverage unpaired human and robot demonstrations to learn visual representations that map both human and robot images into a shared embedding space [4, 6, 13, 14, 15]. A policy is then trained to generate actions conditioned on robot video embeddings, and directly transferred at test time to work with embeddings from a human prompt video. However, a key assumption these works rely on is that the demonstrator and robot perform tasks with matched execution, i.e., the human executes tasks in a visually similar way to that of the robot (e.g. slowly moving one arm with a simple grasp). In reality, humans often act more swiftly, use both hands for manipulation, or even execute multiple tasks simultaneously, creating a mismatch in execution styles. This mismatch leads to misalignment between the demonstrator and robot embeddings, hindering direct policy transfer.
@@
We tackle this problem of imitation under mismatched execution. Our key insight is that while a robot and demonstrator may perform the same task in visually and physically different ways, we can establish a high-level equivalence by reasoning over the entire sequence of image embeddings they generate. We show that while individual image embeddings may appear different between robot and demonstrator, we can construct sequence-level similarity functions where the two are closer. Notably, we can do this without fine-tuning representations on paired data.
@@
We propose RHyME (Retrieval for Hybrid Imitation under Mismatched Execution), a framework which trains a robot policy to follow a long-horizon demonstration from a mismatched expert without access to paired robot-demonstration videos (Fig. 1). First, RHyME defines a sequence-level similarity metric between robot and demonstrator embeddings, using optimal transport to measure alignment. Given a robot trajectory and a database of demonstrator data, RHyME imagines a long-horizon video by retrieving and composing short-horizon snippets of demonstrations similar to the robot video. This retrieval process is guided by an optimal transport similarity metric between robot and demonstrator sequences. The framework trains a policy using a hybrid approach, incorporating both real robot demonstrations and imagined demonstrator sequences. Our contributions are:
@@
    We propose RHyME, a novel retrieval-based algorithm for one-shot initiation from videos of a demonstrator with mismatched execution. Without access to paired datasets, RHyME utilizes high-level equivalence between robot and demonstrator videos using optimal transport costs.
@@
    We systematically study the performance of different video retrieval by proposing three novel cross-embodiment datasets in the Franka Kitchen simulator. The demonstration datasets, totaling 10 hours of videos, are designed to vary in visual and execution mismatches relative to the robot.
@@
    We show that RHyMe outperforms a range of baselines on all degrees of mismatches, yielding over 50% increase in success rates in the most challenging scenarios.
@@ II Related Work
Demonstrations have been utilized to guide robot manipulation policies in several different ways. We place our work relative to each cluster of related research.
@@
Tracking Reference Motion. The imitation challenge is reduced to motion tracking when the robot receives the demonstrator’s motion as input. Robots with human-like joint configurations can directly mimic human trajectories, a method applied to humanoid robots with similar morphologies [16, 17, 18, 19, 20, 21] and robotic hands [22, 23, 24, 25, 26, 27, 28, 29]. When the execution capabilities are mismatched, either human models are simplified [16, 18, 20] or robot trajectories [17, 19, 30, 31] are optimized to approximately match the reference trajectory. [21] shows a method to extract human poses from videos and use that as a reference for training reinforcement learning agents. Similarly, human demonstrations have been used to guide reinforcement learning for robotic hands [22, 23, 24, 25]. More recently, large-scale video datasets of humans on the internet have been used to extract hand positions and adapted for robot manipulation [28, 27, 29, 26, 32], or rely on other common abstractions such as optical flow as a common trajectory representation across embodiments [33], still requiring execution to be matched. Distinct from these works, we focus on learning robot manipulation tasks directly from demonstrators’ RGB videos without explicit motion input.
@@
Learning Reward Functions from Demonstrator Videos. These methods tackle the problem of matching the demonstrator when their motion cannot be simply mimicked. Still, their videos contain useful task information and can be used to learn reward functions for reinforcement learning on the robot. A general method in this line of research [34, 35, 36, 2] is to extract reward functions that encourage the robot to manipulate objects in the same way as the video. For example, [34] self-supervises the robot by making it learn to match the demonstrations of a demonstrator perturbing a rope. GraphIRL [2] extracts sequences of object pose movement from the demonstrator and enforces temporal cyclic consistency with the robot’s execution. Other approaches frame the problem as task matching [37, 38], where the robot is rewarded when it is deemed to perform the same skill as the demonstrator. While these papers tackle the problem of mismatched execution directly, they require reinforcement learning to train robot policies, which can be challenging for complex tasks.
@@
Learning Aligned Robot-Demonstrator Representations. Another strategy for addressing this challenge is to train representations for both the robot and the demonstrator that are indistinguishable when performing the same task. This approach often frames the task as a video translation challenge, where the demonstrator’s video is converted into a robot’s perspective to simplify task mimicry [10, 9, 39]. Additionally, methods like WHIRL [3] align videos by effectively masking out both the robot and the demonstrator, creating a neutral visual field. Embeddings can be aligned using datasets that either directly pair human and demonstrator actions or utilize human preference datasets to rank image frames, as shown in X-IRL [1] and RAPL [40]. Unlike these methods, our approach does not depend on labeled correspondences between robots and demonstrators.
@@
One-shot Visual Imitation from Demonstration Videos. We tackle this problem setting in our work where the robot imitates actions from demonstrator videos in a one-shot setting, i.e., the robot uses a prompt video as a guide, aiming to replicate the demonstrated actions after viewing them once [7, 8, 4, 6, 41, 42, 43, 32]. If a paired dataset of robot and demonstrator videos executing the same task exists, the robot can learn to translate a prompt video into actions directly [7, 8]. The closest to our work is the setting without paired data of demonstrator and robot skills. Prior works [6, 4] train policies conditioning on robot videos and zero-shot transfer to a prompt demonstration at test time using aligned visual embeddings. For example, XSkill [4] uses a self-supervised clustering algorithm based on visual similarity to align representations of demonstrations and robot videos. However, such an approach can falter when there are significant mismatches in execution. We address this issue by posing the visual imitation problem as one of train-time retrieval. During training, we match robot videos to the closest demo videos from an unpaired play dataset to imagine synthetic demo videos. Training robot policies conditioned on these synthetic videos enable the robot to translate demonstration videos into robot actions at test time.
@@ III Problem Formulation
Inference Time: Translate Demonstration Video to Robot Actions. The robot’s goal is to replicate a series of tasks demonstrated in a video using a policy $\pi(a_{R}|s_{R},\mathbf{v_{H}})$ that translates the video into robot actions $a_{R}$ at state $s_{R}$. The demonstration video is a sequence of images $\mathbf{v_{H}}=\{v_{H}^{0},v_{H}^{1},\dots,v_{H}^{T}\}$, where $T$ is the length of the video.
Train Time: Learning from Unpaired Demonstrator and Robot Data. While training a policy with paired robot and demonstrator data is feasible, collecting such data at scale is impractical. Instead, we frame the problem as learning and leveraging aligned embeddings from unpaired data, enabling the transfer of policies trained on robot embeddings to demonstration embeddings.
We assume access to two datasets —a robot dataset ($D_{\rm robot}$) of long-horizon manipulation tasks and a play dataset ($D_{\rm play}$) of short-horizon demonstrator video clips showing interactions with objects and the environment. The robot dataset, $D_{\rm robot}=\{(\mathbf{\xi_{R}},\mathbf{v_{R}})\}$, comprises pairs of state-action trajectories and robot videos. Each robot trajectory, $\mathbf{\xi_{R}}=\{(s_{R}^{0},a_{R}^{0}),(s_{R}^{1},a_{R}^{1}),\dots,(s_{R}^{T},a_{R}^{T})\}$, represents the sequence of robot states and actions throughout an episode. Correspondingly, the video $\mathbf{v_{R}}=\{v_{R}^{0},v_{R}^{1},\dots,v_{R}^{T}\}$ is a sequence of images of the robot executing the task. The play dataset, $D_{\rm play}=\{\mathbf{v_{H}}\}$, consists of demonstrator videos that do not have direct correspondence with the robot dataset. At test time, the demonstrator’s video contains a set of tasks whose composition is unseen by the robot during training. However, we make an assumption that the constituent tasks are individually covered both in $D_{\rm play}$ and $D_{\rm robot}$. This assumption is consistent with prior work [4].
Our goal is to train two modules: a vision encoder and a robot policy. The vision encoder maps both robot and demonstrator videos into a shared embedding space to enable translation. We employ a video encoder $\phi(\mathbf{v})$ to extract a sequence of embeddings $\mathbf{z}=\{z_{0},z_{1},....,z_{T}\}$ for each frame from all videos [^1] We encode a 1-timestep sliding window of 8 neighboring images to generate each image embedding.. Then, given a demonstrator video $\mathbf{v_{H}}$, we generate a sequence of latent embeddings $\mathbf{z_{H}}$. We aim to train a policy that conditions on the sequence of embeddings to predict robot actions $\pi(a_{R}|s_{R},\mathbf{z_{H}})$ without access to paired robot and demonstrator data. We discuss how to train both the encoder and the policy in Section IV.
@@ IV Approach
We present RHyME, a one-shot imitation learning algorithm that translates demonstrations to robot actions, without paired data. Before policy training, we first train a video encoder using a dataset of unpaired robot and demonstration videos (Section  IV-A). Then, this trained video-encoder is frozen and utilized for retrieval during policy training (Section IV-B). At train time, given just a robot trajectory, RHyME imagines a corresponding demonstration by retrieving and composing corresponding demonstrator snippets. It then trains a policy to predict robot actions, conditioned on the imagined demonstration. We discuss details of the retrieval, training process, and video embeddings below.
@@
Input: Robot embeddings $\mathbf{z_{R}}$, demonstrator play dataset $D_{\rm play}$, video encoder $\phi(\mathbf{z}|\mathbf{v})$, segment length $K$, sequence-level video distance function $d$
@@
// Divide sequence into short snippets

### v5

@@ Abstract
Human demonstrations as prompts are a powerful way to program robots to do long-horizon manipulation tasks. However, translating these demonstrations into robot-executable actions presents significant challenges due to execution mismatches in movement styles and physical capabilities. Existing methods either depend on human-robot paired data, which is infeasible to scale, or rely heavily on frame-level visual similarities that often break down in practice. To address these challenges, we propose RHyME, a novel framework that automatically aligns human and robot task executions using optimal transport costs. Given long-horizon robot demonstrations, RHyME synthesizes semantically equivalent human videos by retrieving and composing short-horizon human clips. This approach facilitates effective policy training without the need for paired data. RHyME successfully imitates a range of cross-embodiment demonstrators, both in simulation and with a real human hand, achieving over 50% increase in task success compared to previous methods. We release our code and datasets at this website.
@@ I Introduction
Human demonstrations offer an effective approach for programming robots to execute long-horizon manipulation tasks [1, 2, 3, 4]. Unlike language instructions, demonstrations are grounded in the task environment, providing rich cues for what steps to follow, which objects to interact with, and how to interact with them [5, 6].
@@
We view this as a translation problem where a human video must be translated into a series of robot actions [7, 8, 9, 10]. However, training such policies typically requires paired human-robot demonstrations, which is impractical to collect at scale for long-horizon tasks. Although large-scale human videos (e.g., YouTube) and robot datasets exist [11, 12], they are unpaired, making them unsuitable for directly learning this translation.
@@
Prior works leverage unpaired human and robot demonstrations to learn visual representations that map both human and robot images into a shared embedding space [4, 6, 13, 14, 15]. A policy is then trained to generate actions conditioned on robot video embeddings, and directly transferred at test time to work with embeddings from a human prompt video. However, a key assumption these works rely on is that the human and robot perform tasks with matched execution, i.e., the human executes tasks in a visually similar way to that of the robot (e.g. slowly moving one arm with a simple grasp). In reality, humans often act more swiftly, use both hands for manipulation, or even execute multiple tasks simultaneously, creating a mismatch in execution styles. This mismatch leads to misalignment between the human and robot embeddings, hindering direct policy transfer.
@@
We tackle this problem of imitation under mismatched execution. Our key insight is that while a human and robot may perform the same task in visually and physically different ways, we can establish a high-level equivalence by reasoning over the entire sequence of image embeddings they generate. We show that while individual image embeddings may appear different between human and robot, we can construct sequence-level similarity functions where the two are closer. Notably, we can do this without fine-tuning representations on paired data.
@@
We propose RHyME (Retrieval for Hybrid Imitation under Mismatched Execution), a framework which trains a robot policy to follow a long-horizon demonstration from a mismatched expert without access to paired human-robot videos (Fig. 1). First, RHyME defines a sequence-level similarity metric between human and robot embeddings, using optimal transport to measure alignment. Given a robot trajectory and a database of human play data, RHyME imagines a long-horizon video by retrieving and composing short-horizon snippets of human demonstrations similar to the robot video. This retrieval process is guided by an optimal transport similarity metric between human and robot sequences. The framework trains a policy using a hybrid approach, incorporating both real robot demonstrations and imagined human sequences. Our contributions can be summarized as:
@@
    We propose RHyME, a retrieval-based algorithm for one-shot imitation of long-horizon demonstrations of humans with mismatched execution.
@@
    Without access to paired human-robot datasets, RHyME uses optimal transport to imagine human demonstrations corresponding to each robot trajectory.
@@
    We release three novel datasets for cross-embodiment learning in simulation with increasing levels of execution mismatch compared to the robot. RHyME outperforms baselines in simulation and the real world, improving task completion by $2$x when imitating unseen human videos.
@@ II Related Work
Human demonstrations have been utilized to guide robot manipulation policies in several different ways. We place our work relative to each cluster of related research.
@@
Tracking Reference Motion. The imitation challenge is reduced to motion tracking when the robot receives the demonstrator’s motion as input. Robots with human-like joint configurations can directly mimic human trajectories, a method applied to humanoid robots with similar morphologies [16, 17, 18, 19, 20, 21] and robotic hands [22, 23, 24, 25, 26, 27, 28, 29]. When the execution capabilities are mismatched, either human models are simplified [16, 18, 20] or robot trajectories [17, 19, 30, 31] are optimized to approximately match the reference trajectory. [21] shows a method to extract human poses from videos and use that as a reference for training reinforcement learning agents. Similarly, human demonstrations have been used to guide reinforcement learning for robotic hands [22, 23, 24, 25]. More recently, large-scale video datasets of humans on the internet have been used to extract hand positions and adapted for robot manipulation [28, 27, 29, 26, 32], or rely on other common abstractions such as optical flow as a common trajectory representation across embodiments [33], still requiring execution to be matched. Distinct from these works, we focus on learning robot manipulation tasks directly from human RGB videos without explicit motion input.
@@
Learning Reward Functions from Demonstrator Videos. These methods tackle the problem of matching the demonstrator behaviors when their motion cannot be simply mimicked. Still, their videos contain useful task information and can be used to learn reward functions for reinforcement learning on the robot. A general method in this line of research [34, 35, 36, 2] is to extract reward functions that encourage the robot to manipulate objects in the same way as the video. For example, [34] self-supervises the robot by making it learn to match the demonstrations of a demonstrator perturbing a rope. GraphIRL [2] extracts sequences of object pose movement from the demonstrator and enforces temporal cyclic consistency with the robot’s execution. Other approaches frame the problem as task matching [37, 38], where the robot is rewarded when it is deemed to perform the same skill as the demonstrator. While these papers tackle the problem of mismatched execution directly, they require reinforcement learning to train robot policies, which is challenging for complex tasks.
@@
Learning Aligned Human-Robot Representations. Another strategy for addressing this challenge is to train representations for both the robot and the human that are indistinguishable when performing the same task. This approach often frames the task as a video translation challenge, where the demonstrator’s video is converted into a robot’s perspective to simplify task mimicry [10, 9, 39]. Additionally, methods like WHIRL [3] align videos by effectively masking out both the robot and the human, creating a neutral visual field. Embeddings can be aligned using datasets that either directly pair human and robot actions or utilize human preference datasets to rank image frames, as shown in X-IRL [1] and RAPL [40]. Unlike these methods, our approach does not depend on labeled human-robot correspondences.
@@
One-shot Visual Imitation from Demonstration Videos. We tackle this problem setting in our work where the robot imitates actions from human demonstration videos in a one-shot setting, i.e., the robot uses a prompt video as a guide, aiming to replicate the demonstrated actions after viewing them once [7, 8, 4, 6, 41, 42, 43, 32]. If a paired dataset of human and robot videos executing the same task exists, the robot can learn to translate a prompt video into actions directly [7, 8]. The closest to our work is the setting without paired data of human and robot skills. Prior works [6, 4] train policies conditioning on robot videos and zero-shot transfer to a prompt demonstration at test time using aligned visual embeddings. For example, XSkill [4] uses a self-supervised clustering algorithm based on visual similarity to align representations of human and robot videos. However, such an approach can falter when there are significant mismatches in execution. We address this issue by posing the visual imitation problem as a train-time retrieval problem. During training, we match robot videos to the closest human snippets from an unpaired play dataset to imagine synthetic demonstration videos. Training robot policies conditioned on these synthetic videos enable the robot to translate demonstration videos into robot actions.
@@ III Problem Formulation
Inference Time: Translate Human Demonstration Video to Robot Actions. The robot’s goal is to replicate a series of tasks demonstrated in a video using a policy $\pi(a_{R}|s_{R},\mathbf{v_{H}})$ that translates the video into robot actions $a_{R}$ at state $s_{R}$. The human demonstration video is a sequence of images $\mathbf{v_{H}}=\{v_{H}^{0},v_{H}^{1},\dots,v_{H}^{T}\}$, where $T$ is video length.
Train Time: Learning from Unpaired Human and Robot Data. While training a policy with paired human-robot data is feasible, collecting such data at scale is impractical. Instead, we frame the problem as learning aligned embeddings from unpaired data, enabling the transfer of policies trained on robot embeddings to human embeddings.
We assume access to two datasets —a robot dataset ($D_{\rm robot}$) of long-horizon manipulation tasks and a play dataset ($D_{\rm play}$) of short-horizon human video clips showing interactions with objects and the environment. The robot dataset, $D_{\rm robot}=\{(\mathbf{\xi_{R}},\mathbf{v_{R}})\}$, comprises pairs of state-action trajectories and robot videos. Each robot trajectory, $\mathbf{\xi_{R}}=\{(s_{R}^{0},a_{R}^{0}),(s_{R}^{1},a_{R}^{1}),\dots,(s_{R}^{T},a_{R}^{T})\}$, represents the sequence of robot states and actions throughout an episode. Correspondingly, the video $\mathbf{v_{R}}=\{v_{R}^{0},v_{R}^{1},\dots,v_{R}^{T}\}$ is a sequence of images of the robot executing the task. The play dataset, $D_{\rm play}=\{\mathbf{v_{H}}\}$, consists of human videos that do not have direct correspondences with the robot dataset. At test time, the demonstrator’s video contains a set of tasks whose composition is unseen by the robot during training. However, consistent with prior work [4], we assume that the constituent tasks are individually covered both in $D_{\rm play}$ and $D_{\rm robot}$.
Our goal is to train two modules: a vision encoder and a robot policy. The vision encoder maps both human and robot videos into a shared embedding space to enable translation. We employ a video encoder $\phi(\mathbf{v})$ to extract a sequence of embeddings $\mathbf{z}=\{z_{0},z_{1},\dots,z_{T}\}$ for each frame from all videosWe encode a 1-timestep sliding window of 8 neighboring images to generate each image embedding.. Then, given a human demonstration video $\mathbf{v_{H}}$, we generate a sequence of latent embeddings $\mathbf{z_{H}}$. We aim to train a policy that conditions on the sequence of embeddings to predict robot actions $\pi(a_{R}|s_{R},\mathbf{z_{H}})$ without access to paired human and robot data. We discuss how to train both the encoder and the policy in Section IV.
@@ IV Approach
We present RHyME, a one-shot imitation learning algorithm that translates human videos into robot actions, without paired data. Before policy training, we first train a video encoder using a dataset of unpaired human and robot videos (Section  IV-A). Then, this trained video-encoder is frozen and utilized for retrieval during policy training (Section IV-B). At train time, given just a robot trajectory, RHyME imagines a corresponding demonstration by retrieving and composing short-horizon human snippets. It then trains a policy to predict robot actions, conditioned on the imagined demonstration. We discuss details of the retrieval, training process, and video embeddings below.
IV-A Training the Vision Encoder
We align the human and robot video embeddings in three ways: visually, temporally, and at the task level, all without requiring trajectory-level correspondences. We employ unsupervised losses $\mathcal{L}_{\rm vis}(\phi)$ and $\mathcal{L}_{\rm temp}(\phi)$ for visual and temporal alignment, following prior works [44, 45], and introduce an optional task alignment loss $\mathcal{L}_{\rm task}(\phi)$.
Visual Alignment. To align human and robot embeddings ($\mathbf{z_{R}}$, $\mathbf{z_{H}}$), we use SwAV [44], a self-supervised method that clusters images based on shared visual features. SwAV learns a set of $K$ prototype vectors, to which each image is assigned. The SwAV loss $\mathcal{L}_{\rm vis}(\phi)$ updates both the encoder and prototypes, aligning human and robot videos by clustering similar visual features.
Temporal Alignment. To align temporally adjacent frames in human and robot videos, we use Time Contrastive Loss [45]. This loss encourages embeddings of frames close in time to be similar. For each frame $z^{t}$, we define a positive set $\mathbf{z^{+}}$ of frames within a temporal window $w$, and a negative set $\mathbf{z^{-}}$ for frames outside this window. Using the contrastive loss $\mathcal{L}_{\rm temp}(\phi)$, we pull embeddings from the positive set closer and push negative set embeddings further apart, capturing temporal continuity across videos.
Task Alignment. Task-level alignment $\mathcal{L}_{\rm task}(\phi)$ is used when a small set of paired human and robot snippets is available. Unlike frame-level methods, this aligns video embeddings of the robot $\mathbf{z_{R}}$ and demonstrator $\mathbf{z_{H}}$. We compute the optimal transport distance $d(\mathbf{z_{R}},\mathbf{z_{H}})$ to measure the similarity between two sequences of video embeddings. We then apply a contrastive learning objective (INFO-NCE [46]) to pull matched embeddings closer and push different-task embeddings apart. The final task alignment loss is:
$\mathcal{L}_{\rm task}(\phi)=-\sum\limits_{i}\frac{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))}{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))+\sum_{j\neq i}\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{j}))}$
Our final loss function for training the visual encoder $\phi$ is:
$\mathcal{L}(\phi)=\lambda_{\rm vis}\mathcal{L}_{\rm vis}(\phi)+\lambda_{\rm temp}\mathcal{L}_{\rm temp}(\phi)+\lambda_{\rm task}\mathcal{L}_{\rm task}(\phi)\vspace{-2mm}$ (1)
where $\lambda_{\rm task}=0$ by default and non-zero only with access to short-horizon paired data.
IV-B Training the Robot Policy
Training Overview. Algorithm 1 details our approach to train robot policy $\pi_{\theta}$ using both robot trajectories and imagined human demonstration videos. The training process has two stages.
@@ IV-B Training the Robot Policy
Input: Robot Dataset ${D}_{\rm robot}$, Human Play Dataset $D_{\rm play}$, Video Encoder $\phi(\mathbf{z}|\mathbf{v})$
Output: Trained Robot Policy $\pi_{\theta}(a|s,\mathbf{z})$
Initialize Robot Policy $\pi_{\theta}$
1while not converged do
Get robot video and actions $\mathbf{\xi_{R}},\mathbf{v_{R}}\sim D_{\rm robot}$
Generate robot embeddings $\mathbf{z_{R}}=\phi(v_{R})$
// Retrieve human embeddings
// Hybrid Training
2      for $(s_{t},a_{t})$ in $\xi_{R}$ do
// Condition on imagined demo
[…]
// Condition on robot video
[…]
4
Return Trained Robot Policy $\pi$
Algorithm 1 RHyME: Retrieval for Hybrid imitation under Mismatched Execution
Stage 1: Create a Paired Dataset. For each robot trajectory $\mathbf{\xi_{R}}$ and video $\mathbf{v_{R}}$ in $D_{\rm robot}$, we encode the robot video into embeddings $\mathbf{z_{R}}=\phi(\mathbf{v_{R}})$ using the learned video encoder $\phi$. We then retrieve imagined human embeddings $\mathbf{\hat{z}_{H}}$ by aligning $\mathbf{z_{R}}$ with demonstration snippets from the play dataset $D_{\rm play}$, through the function Imagine-Demo. This produces a paired dataset $D_{\rm paired}$ containing $(\mathbf{\hat{z}_{H}},\mathbf{z_{R}},\mathbf{\xi_{R}})$.
Stage 2: Train Policy on Paired Dataset. The policy $\pi_{\theta}$ is trained on the paired dataset $D_{\rm paired}$ in a hybrid fashion. For each element in $D_{\rm paired}$, we update the policy in two modes —Mode 1: The policy is conditioned on the robot video embeddings $\mathbf{z_{R}}$ to predict actions $\pi_{\theta}(a_{t}|s_{t},\mathbf{z_{R}})$, Mode 2: The policy is conditioned on the imagined human demonstration embeddings $\mathbf{\hat{z}_{H}}$ to predict actions $\pi_{\theta}(a_{t}|s_{t},\mathbf{\hat{z}_{H}})$. By alternating between these two modes, the policy learns to generalize from both robot and imagined human videos, enabling it to handle execution mismatches.
3
Input: Robot Embeddings $\mathbf{z_{R}}$, Human Play Dataset $D_{\rm play}$, Video Encoder $\phi(\mathbf{z}|\mathbf{v})$, Segment Length $K$, Distance Function $d$
@@
// Divide long-horizon robot sequence into short-horizon clips